### v3

# Intent-Based Access Control: Using LLMs to Intelligently Manage Access Control

Abstract.
In every enterprise database, administrators must define an access control policy that specifies which users have access to which assets. Access control straddles two worlds: policy (organization-level principles that define who “should” have access) and process (database-level primitives that actually implement the policy). Assessing and enforcing process compliance with a policy is a manual and ad-hoc task. This paper introduces a new paradigm for access control called Intent-Based Access Control for Databases (IBAC-DB). In IBAC-DB, access control policies are expressed more precisely using a novel format, the natural language access control matrix (NLACM). Database access control primitives are synthesized automatically from these NLACMs. These primitives can be used to generate new DB configurations and/or evaluate existing ones. This paper presents a reference architecture for an IBAC-DB interface, an initial implementation for PostgreSQL (which we call LLM4AC), and initial benchmarks that evaluate the accuracy and scope of such a system. We further describe how to extend LLM4AC to handle other types of database deployment requirements, including temporal constraints and role hierarchies. We propose RHieSys, a requirement-specific method of extending LLM4AC, and DePLOI, a generalized method of extending LLM4AC. We find that our chosen implementation, LLM4AC, vastly outperforms other baselines, achieving high accuracies and F1 scores on our initial Dr. Spider benchmark. On all systems, we find overall high performance on expanded benchmarks, which include state-of-the-art NL2SQL data requiring external knowledge, and real-world role hierarchies from the Amazon Access dataset.
1. Introduction
Over the last several years, there has been a proliferation of easy-to-use software tools for data science. Organizations of all sizes and from all industries are applying data-driven models to make decisions. This trend, which is often called the democratization of data science (Chou et al., 2014), is as worrying as it is exciting. On one hand, some of humanity’s most difficult scientific challenges are being addressed with data science from particle physics to drug discovery. On the other hand, there are numerous cautionary tales of data misuse, data breaches, data quality issues, and improper data sharing (Monroe-White et al., 2021; Janssen et al., 2020; Kerber, 2020; Seh et al., 2020; Fahey, 2021). Configuring database deployments, to manage availability, usability, integrity, and security of the data in enterprise systems, has consequently emerged as a key discipline in enterprise data management. Access control is one of the most important technical challenges when configuring databases, where a system must ensure that only authorized users can access particular data assets.
In any sizeable organization, it is impractical to exhaustively map individual users to data assets. Thus, role-based access control (RBAC) was a key innovation and is widely supported by data management platforms where groups of users can be selectively given access (Sandhu et al., [n.d.]). Such forms of access control are well-studied in enterprise SQL databases, where roles, views, and GRANT statements are already used to restrict user access to data (Chaudhuri et al., 2011; Bertino et al., 2011; Bertino and Sandhu, 2005). Similarly, many databases additionally support discretionary access control (DAC) where roles can be given delegation privileges. RBAC and DAC help reduce the administrative burden in defining access control rules for a database.
Despite these features, access misconfiguration is still widely described as a major cyber-security risk and a business cost (Son et al., 2013). Access control fundamentally straddles two different worlds: policy and process. Policy describes an organization’s intentions with respect to access control, data retention, authorization, etc. These are broad principles that govern how an organization chooses to delegate access to data assets. Policies are often constructed with legal or regulatory advice and are specified in documents, reports, and guidelines. Process, on the other hand, describes the translation of these policies into actionable code and rules that implement such functionality. Going from policy to process often requires identifying which text in a document corresponds to access control rules in a database, and which database components (schema elements, database roles, etc.) these rules reference.
Figure 1. Access Control Policies (Written in Natural Language Documents) to Process (Written in SQL GRANT statements)
For example, consider the access control policy document shown in Figure 1. These documents form a critical part of an organization’s security infrastructure and are needed for compliance (aic, [n.d.]). An access control policy document is typically written in natural language (NL), and often contains statements that are not relevant for database access control (e.g., job descriptions, such as “Marketing managers are responsible for making decisions based on customer data and product sales.”). Further, even after one has identified the relevant parts of a policy document (e.g., “Marketing managers can look at customer data. They are responsible for granting their read privileges to regional managers, for regional customer data.”), one must identify which permissions are being given on which tables/views for which roles/users, and generate the appropriate SQL role/user creation statements, view creation statements, and GRANT statements. Lastly, once an access control policy has been implemented in a database, one should be able to audit the implementation to determine whether the policy was implemented correctly.
Often, the individuals or groups in charge of developing policy, e.g., security experts, organizational counsel, and insurance providers, do not have the expertise to advise their correct implementation in a database system. As the database ecosystem grows with new tools, each one has its own access control idiosyncrasies and specification languages. Today’s IT practice leaves both policy makers and database administrators with a lack of information that would help verify that a given access control implementation matches an organization’s intent. As a step towards a solution, this paper studies directly specifying access control policies in natural language, and automatically implementing and auditing such policy implementations for correctness. Specifically, we make the following contributions:
  • IBAC-DB: A new paradigm for database access control called Intent-Based Access Control for Databases (IBAC-DB) [^1] Inspired by similarly named techniques in network and firewall configurations.. An IBAC-DB system does not just specify access control rules, but can intelligently evaluate these rules against natural language policies to identify potential violations.
  • LLM4AC for Access Control Privileges: A system built from a reference architecture for an IBAC-DB implementation that provides a novel format for specifying access control policies, automatically implementing them, and comparing implementations to policies for correctness.
  • Differencing Definition and Algorithm: A novel definition and algorithm for differencing access control policies with their implementations to compare policies with their implementations.
  • Generalizing to Other Deployment Requirements: we propose two other systems that generalize LLM4AC to other requirements needed to deploy databases: one requirement-specific system that synthesizes and differences implementation using LLM prompts specialized to multiple types of deployment requirements, such as other access control requirements (e.g., role hierarchies, temporal access control), and one generalized system that uses user-inputted requirement types to analyze the access control policy for these requirement types, and synthesize and difference the corresponding access control implementation.
  • Differencing Benchmark Design: a novel benchmark for testing that an IBAC-DB system can correctly determine whether an access control policy matches its implementation.
Specifically, IBAC-DB enables easy identification and extraction of database access control rules from NL using the natural language access control matrix (NLACM), a novel abstraction. We implement access control policies by adapting NL2SQL systems to generate access control SQL, a capability which no current NL2SQL systems possess. We compare a policy to its implementation using a novel differencing procedure, which uses LLMs to determine which access control rules were implemented that do not comply with the policy. We find that the accuracy of this procedure increases greatly when we prune the candidate matches input to the LLM that do not have matching DB literals first. This allows LLM4AC to achieve near-perfect F1 scores on our benchmarks (e.g., F1 scores of 0.99).
This paper is structured as follows: in Section 2, we describe the key definitions, problem statement, and related work describing synthesis and auditing of access control. In Section 3, we describe a reference architecture specifically for implementing access control privileges (permitted SQL operators to roles on database views). In Section 4, we describe our system, LLM4AC, which is an instantiation of the reference architecture using ChatGPT and Postgres. In Section 5, we evaluate LLM4AC. In Sections 6 and 7, we describe and evaluate our requirement-specific system design respectively. Similarly, in Sections 8 and 9, we describe and evaluate our generalized system design. We conclude in Section 10.
2. Background
First, we introduce some notation and formalism to scope the IBAC-DB problem.
2.1. Access Control Definitions
Let us consider access control to a single database. Over this database, there are a set of possible users $U$. These users can run a set of possible SQL statements $S$ over this database. An access control scheme is a set of rules that specifies who should be given access under what conditions. Formally, an access control scheme for a database is a program, that determines whether a particular user is allowed to run the desired SQL statement over the database:
$\mathcal{P}(user,stmt):U\times S\mapsto\{allowed,denied\}$
$\mathcal{P}$ describes a literal implementation of access control for a particular database. For example, in an SQL database, $\mathcal{P}$ is defined as a sequence of GRANT statements. However, such programs are not created in a vacuum. An access control policy is a set of rules that specifies who should be given access to what resources. For example, an organization’s regulatory posture may dictate who has access to data (e.g., only relevant doctors may view electronic health records). Or, an organization’s security team may implement a particular minimal privilege strategy (e.g., only senior management can modify customer records). Whatever the reason, these decisions are described at a higher level of abstraction than $\mathcal{P}$.
SOC-2 compliance requires that access control policies be specified in an access control matrix (aic, [n.d.]). An access control matrix defines rows that are users (or groups) and columns that are data assets. The cells of this matrix specify the allowed operations of that user on the asset. Similarly, research institutions that work with identifiable information need to submit data management plans written in document form to institutional review boards and funding agencies (loy, 2020; nwp, 2022). In general, access control policies are written in natural language or quasi-natural language (e.g., an access control matrix) by a database non-expert. To correctly implement such policies, one must determine which parts of the policy represent access control rules for the database, and which components of the database those parts refer to. This can require multiple rounds of communication between a policy expert and a database expert (e.g., a database administrator) (Eryurek et al., 2021; nis, [n.d.]; sat, 2023; act, [n.d.]).
2.2. Auditable Access Control
We have to compare two access control schemes $\mathcal{P}$, which is a scheme specified on the database in SQL, and $\mathcal{P}^{*}$, which is a scheme specified in natural language. Hereafter, we will refer to $\mathcal{P}$ as an access control “implementation” and $\mathcal{P}^{*}$ as an access control “policy”. We fully understand that there is inherent imprecision in natural language and there might not exist an unambiguous $\mathcal{P}^{*}$; however, we scope our initial exploration of this problem in such a way to avoid such ambiguities. The main goal is to be able to check for compliance, or formally:
Definition 0 (Compliance).
An access control implementation $\mathcal{P}$ is compliant against a policy $\mathcal{P}^{*}$ if and only if
${}~{}\forall(u,t)\in U\times T:~{}~{}\mathcal{P}(u,t)\implies\mathcal{P}^{*}(u,t)$
This definition states that every allowed SQL statement in $\mathcal{P}$ is also allowed in $\mathcal{P}^{*}$, or in other words, $\mathcal{P}$ is at least as restrictive $\mathcal{P}^{*}$. Many solutions have been developed that enable automated implementation and verification of access control policies. However, we will see that these come at the cost of restricting access control rules that can be expressed.
One solution is to adhere to an access control paradigm for specifying access control policies, such as role-based access control (RBAC), discretionary access control (DAC), etc. Adhering to these paradigms alone can greatly simplify defining and implementing access control policies. For example, if one defines a role-based access control policy on a database, one need only implement it by concatenating the SQL describing the roles, tables/views, and privileges into a SQL GRANT statement. However, implementing such policies through a database system alone requires database expertise and explicit enumeration of the complete set of access control rules for all database users.
To reduce the expertise and manual effort needed to implement an access control policy, prior work has focused on allowing access control to be defined via programming languages for access control (Hu Vincent, [n.d.]; Zigmond et al., 2019; xac, [n.d.]), which can then be translated into database privileges, often with guarantees on the correctness of translation, or automated verification via model-checking.
These programming languages allow automatic and provably correct implementation of access control policy. However, they restrict the access control rules that can be expressed, compared to natural language. For example, ShillDB, a recent contract language for database access control, cannot express access control rules on nested queries (Zigmond et al., 2019), but natural language could easily express access control rules on such a query (e.g., Q4, the order priority checking query, from TPC-H).
Based on existing access control solutions, we find that on the one hand, one can use NL to represent policies, but then implementation and auditing for correctness are manual. On the other hand, you can use access control paradigms or languages which automatically and correctly implement access control policies, but then the expression of access control rules is limited.
2.3. Problem Statement
Therefore, in this paper, we solve the problem: how do we enable access control policies to be expressed in natural language, while still allowing automated, correct implementation and auditing of access control schemes? We address this problem by making the following contributions:
  1. (1)
    We propose a new paradigm for access control called Intent-based Access Control for Databases (IBAC-DB), which uses the natural language access control matrix (NLACM) as the input.
  2. (2)
    We propose LLM4AC, a reference architecture for an IBAC-DB interface.
  3. (3)
    To allow for policy auditing, we define differencing, a procedure for comparing an access control policy and its implementation. Note that we cannot formally verify the correctness of implementations, due to the imprecision of natural language. However, we find that our differencing procedure for auditing policies is mostly correct in practice.
2.4. Existing NL2SQL Capabilities
NL2SQL systems seem helpful in implementing NL access control policies as SQL. Current NL2SQL capabilities focus on converting natural language questions about information in a database to queries (typically SELECT queries) (Dong et al., 2023; Yu et al., 2018; Ni et al., 2023; Wang et al., 2020). The state-of-the-art NL2SQL methods develop procedures for prompting LLMs to generate accurate SQL, such as ChatGPT (Dong et al., 2023) or GPT-4 (Gao et al., 2023).
However, to our knowledge, current NL2SQL benchmarks do not contain examples that generate access control queries (Yu et al., 2018; Zhong et al., 2017; Sen et al., 2020), and the database system used for evaluation in most cases is SQLite, which does not support access control.
Further, it is not obvious how to bridge this gap between state-of-the-art NL2SQL methods and the access control use case. For example, enhancing LLM-backed NL2SQL systems by adding naive prompting of ChatGPT can produce an incorrect answer. When given the prompt, Write the postgresql commands to implement such access control, given the database schema, two of the queries generated by ChatGPT are ”GRANT CREATE ON ALL TABLES IN SCHEMA public TO data_architect;” and ”GRANT CREATE ON ALL VIEWS IN SCHEMA public TO data_architect;”. The second query is unnecessary, and uses the word ”VIEWS” in the grant statement which does not exist.
Simply using NL2SQL methods as-is by inputting access control policy documents can also be inaccurate. Consider a straightforward access control rule: ”Grant the user John select access on the customer table with the option of passing down this privilege.” C3 (Dong et al., 2023) translates this NL to the SQL ”GRANT SELECT ON customer TO John;”. This is incorrect because it ignores the option of passing down the privilege–the query must be suffixed with a ”WITH GRANT OPTION”.
Figure 2. Comparison of NL2SQL Translation by LLM method, with respect to syntax and Forward-Reverse-Forward (FRF) Accuracy. Columns are access control policies.
On the other hand, we find that C3 more accurately generates all other GRANT statements when given a natural language sentence that clearly spells out the role, table/view, and privilege, compared to naive prompting, or prompts engineered using typical prompt engineering techniques (e.g., few-shot learning, chain-of-thought prompting, etc.). In Figure 2, we compare naive prompting of ChatGPT (the ”Naive” method), a chain-of-thought prompting solution (the ”Rolled” method), and C3, a state-of-the-art NL2SQL system, with respect to accuracy of the generated syntax, and forward-reverse-forward (FRF) accuracy. We systematically perform this comparison on access control policies containing role hierarchies of various complexities, view creation, and discretionary access control. We synthetically generate 100 policy documents of each type, and compute the accuracy based on the syntax/FRF accuracy of the synthesized SQL script. We see that C3’s syntax and FRF accuracy far exceeds prompting solutions. This suggests that altering NL2SQL methods for access control may facilitate implementation of access control in databases.
Based on these examples and initial experiments, rather than trying to train NL2SQL systems, we hypothesize if access control policies can be formatted to clearly reflect a role, privilege, and view, and we can correctly represent access control policies in terms of the database roles, views, and privileges that correspond to the policy, then NL2SQL methods adapted for access control can correctly implement access control policies.
Why Trust LLMs for Access Control? There are salient differences between the traditional NL2SQL problem for generating SQL DML, and the NL2SQL problem for DDL (including access control). NL ambiguity for current NL2SQL has been observed to lead to multiple correct answers (Floratou et al., 2024) but access control has only one correct implementation (only one set of GRANT statements can correctly implement a database access control policy). Considering this, and the very limited (ideally non-existent) scope for error in access control implementation, why should we trust LLMs for access control? In this paper, we do not assume the output of a NL2DDL system will directly be used by an organization. Instead, we propose a system that will make access control implementation less tedious and error-prone if used in the following workflow: (1) Users will write access control policies (2) The access control policies will be inputted to the system (3) The system will synthesize SQL, and audit the synthesized implementation with respect to the policy (4) The end user will receive the implementation and audit report, which they will use to correct implementation errors. The ease and reliability with which the end user can correct the output implementation will depend on the system’s synthesis and auditing accuracy.
This workflow requires accurate methods of evaluating NL2SQL systems for DDL. NL2SQL systems for DML are evaluated via execution accuracy by comparing outputted recordsets to one or more ground truths. On the other hand, executing access control SQL only changes the database state, and returns nothing. Therefore, a NL2SQL system for access control would currently have to be evaluated by: (i) manually devising the only ground truth implementation, and then (ii) combinatorially generating tests to compare whether allowed operations under the ground truth are allowed under the system-generated implementation, and vice versa (Bertolino et al., 2018; Li et al., 2015; Zhang et al., 2023). These steps are tedious and error-prone. This is why we propose auditing implementation by using LLMs to directly compare the NL policy to its SQL implementation.
3. Access Control via Intents
In this section, we explain what an access control policy intent is, and bridge the gap between NL policy and process for RDBMSs via a new access control scheme called Intent-based access control for databases, or IBAC-DB. Then, we present our system for enabling this paradigm, LLM4AC.
3.1. Specifying Intents
Database access control rules are specified as GRANT statements, each of which consists of the allowed SQL operators for a role/user on a table/view. We refer to NL or SQL that references key elements of access control rules (roles/users, tables/views, or permitted SQL operations) as intents. Then, intent-based access control (IBAC-DB) enables the specification of access control policies using intents. IBAC-DB achieves this using a novel abstraction called the natural language access control matrix, or NLACM. A NLACM specifies database access control rules as follows: a row represents privileges for a database role/user, a column represents privileges for a table/view, and a cell represents the allowed SQL operators. In this way, each matrix cell represents an access control rule.
Definition of NLACM.: Let $D$ be a database schema, which consists of: (i) table schema definitions, (ii) view definitions, (iii) role and user definitions. Then, a NLACM is a $m\times n$ matrix where each row represents the database privileges of a role/user in the database, and each column represents a table/view. The $(i,j)$th cell represent the privileges of role $i$ on table/view $j$.
NLACMs specify access control rules not already implemented in the database. Toward this, they have the following constraints:
  • Each role/user appears in only one row of the NLACM.
  • Each view appears in only one column of the NLACM.
  • (Principle of Failsafe Defaults): There can be roles, views in the DB that do not appear in the NLACM, and we assume they do NOT have any privileges defined.
  • Empty cells are permitted, indicating no privilege is assigned.
  • Non-empty cells indicate which SQL operators are permitted for a given role/user on a given view.
  • Each role and each view can be expressed either as NL or SQL, and the privileges can be a list of SQL operators, or NL.
  • Columns are table names expressed in NL, or view definitions that appear in the database.
We provide a solution for specifying database access control policies in NL whose implementation is in SQL, but our solution is compatible with any language for implementing database access control. This is because of how database access control is implemented in RDBMSs. Whenever a user issues SQL queries, the DBMS uses specialized algorithms to check the query against GRANT statements (Bertino et al., 1996, 2000). Therefore, access control is carried out according to the SQL standard rather than an implementation involving system internals.
3.2. The Benefits of NLACMs
Using NLACMs for access control policies instead of plain NL has three key benefits: (i) NL access control policy documents often contain information irrelevant to implementing the policy in a database. The structure of NLACMs guarantee that all NL will be relevant to the database. (ii) NL is often ambiguous with respect to the access control rules. The structure of NLACMs alleviate this ambiguity. (iii) NLACMs maintain much of the expressivity of NL for access control despite their structure. In this section we discuss these benefits of NLACMs in depth.
NL policies often contain irrelevant information to implementing access control policies. For example, an access control policy template (act, [n.d.]) has a sentence, This policy is the property of CompanyName and is intended for internal use only. It should not be reproduced, partially or wholly, in whatever form.. Another example is, A database administrator is responsible for the usage, accuracy, efficiency, security, maintenance, administration and development of an organization’s computerized database(s), providing support to some or all departments depending on the size of the organization.. These sentences are clearly not intended to be implemented as GRANT statements on a database. On the other hand, the NLACM’s structure lends itself precisely to database access control rules: the $(i,j)$th cell represents the allowed SQL operators for role/user $i$ on table/view $j$.
Even when NL describes database access control rules, it can be very ambiguous. For example, one database access control policy says, SELECT privilege grants a user’s access on views and tables should be limited to authorized personnel.. It is unclear whether this sentence is intended as a general guideline, or a sentence that should result in GRANT statements on the database. Specifically, it is unclear whether to assume that authorized personnel have SELECT privileges on all tables and views in the database. It is also unclear which roles and users are being referenced by ”authorized personnel”. If this sentence appeared in a NLACM, it would make it clear that this sentence is not a general guideline, and there should be a one-to-one or one-to-many relationship between ”authorized personnel” and the database roles/users.
Although NLACMs are a constrained version of NL, they can still be used to express many access control constraints using NL. For example, NLACMs can represent role hierarchies (e.g., a NLACM entry says, ”Same as role x”), subsumed views (e.g., a NLACM column name says, ”the first 100 rows of View 1”), and dynamic access control constraints (e.g., ”read access only from 9am-5pm on weekdays”). Lastly, NLACMs allow IBAC-DB to subsume RBAC (role-based AC) and DAC (discretionary AC), because IBAC-DB intents can be written in NLACMs using plain SQL. In this case, the NLACM becomes a standard ACM representing RBAC and DAC policies.
Figure 3. The LLM4AC Architecture
3.3. A Reference Architecture for IBAC-DB
We propose a reference architecture and system, LLM4AC, for an IBAC-DB interface (see Figure 3 for the full architecture). LLM4AC synthesizes access control policies and compares for the compliance of the implementation as follows: it takes a NLACM expressing an access control policy and a DB schema as input, synthesizes database access control rules (GRANT statements), and then assesses the compliance of the implementation by generating a NLACM from the DB access control rules and differencing the resulting NLACM and the NLACM expressing the policy. We explain each of these steps in detail below and explain the design space for a system built using this reference architecture.
Figure 4. LLM4AC Synthesizer Procedure for Policy Implementation
Synthesizer. The synthesizer procedure is shown in Figure 4. LLM4AC takes a NLACM representing a policy and a DB schema as input. Then, we use a LLM-backed NL2SQL system, such as C3(Dong et al., 2023), which we modify for access control by implementing the novel prompting strategies for generating tables/views, privileges, and GRANT statements from NL in Figure 4. Specifically, we add rules for generating access control SQL and accompanying few-shot examples. We also add chain-of-thought prompting to generate privileges from NL descriptions. The output is a SQL file consisting of GRANT statements, and CREATE ROLE/USER statements.
Generator. The generator procedure simply extracts the list of role/user-privilege-table/view triples from the database (either in the form of a list of GRANT statements issued, or a metadata table containing privileges), and constructs an ACM from these triples.
Figure 5. LLM4AC Differencer Procedure for Policy Auditing.
Differencer. The differencing procedure is shown in Figure 5. The differencer takes two NLACMs as input and determines if the second NLACM is in violation compared to the first. That is, either there are roles or views that are assigned privileges in the second NLACM that are not in the first NLACM, or the same role or view in the second NLACM has more permissive privileges than in the first. The differencer performs the following high-level steps: (a) label each role, view, or privilege in the NLACM as NL or SQL, (b) determine which role-view pairs are shared by both NLACMs by first identifying role-view pairs that share literals from the database, and then giving only these candidate pairs to LLM prompts specialized for comparing NL vs NL and NL vs SQL, (c) for the role-view pairs deemed identical by the LLM, determine whether the privileges are subsumed. The output of differencing will be the role-view pairs and privileges that are in violation.
The Design Space. (1) DBMS SQL Syntax Rules Different DBMSs have varying support for access control. For example, SQLite’s SQL variant has no constructs for access control at all, while PostgresQL supports table and row-level access control, but must implement column-level access control indirectly. Databricks SQL has SQL keywords specifically for column-level access control. Each access control SQL variant requires different rules and few-shot examples for the synthesizer to be successful. (2) LLM Choice Different LLMs are trained to generate different variants of SQL with varying levels of accuracy. Therefore, the LLM must be chosen to align with the desired SQL variant. (3) NL2SQL System Choice Similarly to (2), there are many choices for NL2SQL systems. We recommend (and use) LLM-backed systems, as they are more easily adaptable for generating many variants of SQL, using rules/few-shot prompting, instead of more expensive fine-tuning.
3.4. LLM4AC: Our System
In this paper, we build a system based on our reference architecture, LLM4AC. LLM4AC uses ChatGPT as the LLM, and PostgresQL as the database. slightly alter C3’s existing prompt for translating NL to SQL to also translate view creation, we use chain-of-thought prompting (Wei et al., 2023) to generate privileges from NL descriptions.
4. LLM4AC, In Depth
4.1. The Synthesizer, In-Depth
Definition 0 (The Synthesizer Problem).
Given a NLACM $M$ and a database $D$, consisting of table/view definitions, roles, and users, generate the set of GRANT statements over $D$ implementing the access control rule specified in each cell of $M$.
Prior Work. As mentioned in Section 2, the closest prior work for synthesizing SQL GRANT statements from NL is NL2SQL systems. There are many state-of-the-art NL2SQL systems. Some of these systems train separate models using grammars to avoid syntax errors (Fu et al., 2023). Many of them leverage LLMs (e.g., by fine-tuning LLMs and/or using prompt engineering, including few-shot prompting and describing rules for translation). However, none of these systems are built to generate GRANT or DDL statements.
To bridge this gap, we observe that because many LLM-backed state-of-the-art NL2SQL systems use fine-tuning and prompt engineering, adapting them for access control may simply require adding prompts or examples specific to access control applications.
We enhance one such system which has been observed to be state-of-the-art (Yu et al., 2018), C3 (Dong et al., 2023). C3 is a prompt engineering-based NL2SQL system that uses few-shot examples that include rules to avoid specific syntax errors, and completion prompting to boost performance. We enhance C3 with new access control-specific rules, few-shot examples, and prompts that allow it to accurately translate NL for GRANT statements. We call our variant of C3 C3-AC.
We now describe the synthesizer procedure in detail. Given a NLACM and a DB with no roles or privileges,
  1. (1)
    Extract the roles/user strings from the NLACM and generate a ”CREATE ROLE/USER…” statement for each role/user.
  2. (2)
    Extract the table/view SQL from the NLACM. generate a ”CREATE VIEW…” statement for each table/view in the NLACM. We achieve this using the following prompt: Complete postgres SQL statement only and with no explanation, and do not grant privileges on tables, roles, and users that are not explicitly requested in the statement. CREATE VIEW. This completion-style prompt is similar to that used by C3, but we have adapted it for access control.
  3. (3)
    Map the privileges to SQL operators. Once we have mapped the NLACM roles, tables, and views to the DB, we then map the natural language privileges of the resulting Type 1 NLACM (the roles and views are found directly in the database, but the privileges can be expressed in natural language). We use the following prompt: Consider the following statement: ¡NL for privileges¿. According to this, which of the database operations SELECT, UPDATE/INSERT, DELETE, CREATE, GRANT are permitted for role ¡r¿ on table/view ¡v¿?
  4. (4)
    Synthesize a SQL script from the CREATE ROLE/USER statements, CREATE VIEW statements, and GRANT statements synthesized at each step.
LLMs are known to have syntax errors when generating SQL (Fu et al., 2023). We observe that C3 is usually imprecise in one specific way: when DB literals are used in the NL to be implemented, C3 may generate these literals incorrectly when synthesizing SQL for views (e.g., ”Live_Final” becomes ”Live Final”, ”More then 80 centimeters” becomes ”¿80cm”, etc.). When such queries fail to execute, we repair them using the following procedure:
  1. (1)
    Parse the query for its DB literals (and exclude SQL keywords), column names, and table names.
  2. (2)
    Use the table and column names to query the specific columns used by the query for values contained in the NL. Call the resulting list of DB literals $R$.
  3. (3)
    Embed the literals $R$ using BERT.
  4. (4)
    For each parsed literal from the SQL, use nearest-neighbor search on $R$ to find the highest-matching literal.
  5. (5)
    Replace the literal in the SQL query with the top match from $R$.
4.2. The Differencer, In-Depth
A key problem when implementing access control policies is auditing the implementation for compliance. We cannot verify compliance without a deterministic method for comparing NL and SQL, but we can use LLMs to audit the implementation for compliance. We will show empirically that our policy auditing procedure is mostly accurate. Concretely, we audit policy implementations for compliance by comparing the NLACM containing the NL that expresses policy, and the ACM constructed from privileges implemented on DB. We compare two different policies by comparing two NLACMs that are meant to express policies.
Considering this goal, we define differencing according to the following intuition: given NLACM 1 and NLACM 2, where NLACM 1 represents an access control policy, and NLACM 2 represents its implementation on the database, we want to know whether NLACM 2 matches NLACM 1, and if not, then does NLACM 2 at least grant fewer privileges than NLACM 1. Considering this, we define the difference between NLACM 1 and NLACM 2 as: (i) the roles present in NLACM 2 that are not in NLACM 1, (ii) the views present in NLACM 2 that are not in NLACM 1, (iii) the privileges that are more permissive on the same role-view pair in NLACM 2 than on NLACM 1. By this definition, if there is a difference between NLACM 1 and NLACM 2, we can also conclude that NLACM 2 is in violation with respect to NLACM 1.
Formally, we define the difference between two NLACMs as follows:
Strict Inequality (LLM4AC’s default).: Which database roles are in $M_{2}$, but not in $M_{1}$? Which database tables/views are in $M_{2}$, but not $M_{1}$? Because we assume the principle of failsafe privileges, any role/view in $M_{2}$ but not $M_{1}$ implies that $M_{2}$ is granting privileges on roles/views that $M_{1}$ is not, meaning $M_{2}$ is in violation compared to $M_{1}$. Lastly, for each role $r$ and view $v$ shared between $M_{1}$ and $M_{2}$, are the privileges assigned to $r$ on $v$ in $M_{2}$ subsumed by those of $M_{1}$? If not, then $M_{2}$ is in violation.
Full Procedure. We now write out the full procedure, both steps and prompts, below.
Concretely, given NLACMs $M_{1}$ and $M_{2}$, let $R_{1}$ be the roles of $M_{1}$ and $R_{2}$ be the roles of $M_{2}$. Similarly, let $V_{1}$ be the views of $M_{1}$ and $V_{2}$ be the views of $M_{2}$. Let $R_{1N}$ be the set of roles expressed in natural language, and $R_{1D}$ be the set of roles expressed in SQL. Similarly, let $V_{1N}$ be the set of views expressed in natural language, and $V_{1D}$ be the set of views expressed in SQL. Assume similar naming conventions for $M_{2}$. Then, we compare $M_{1}$ and $M_{2}$ by choosing a role/view from all roles/views in $M_{2}$ each role/view of $M_{1}$ (i.e., $\forall r\in R_{1}$, we match $r$ to a role in $R_{2}$).
We use different prompts for the cases: (i) $r$ is NL, and we compare to natural language roles, $R_{2N}$. (ii) $r$ is NL, and we compare to SQL roles $R_{2D}$, (iii) $r$ is SQL, and we compare to NL roles $R_{2N}$. We similarly use different prompts for these cases when dealing with comparing views.
Then, the full procedure is:
  1. (1)
    Classify the roles/views of each of $M_{1}$ and $M_{2}$ as NL or SQL by searching for SQL keywords in the text.
  2. (2)
    Determine which roles and views are the same, and which are different using a prompt for NL roles/views to NL roles/views, or the NL to SQL prompt chosen to lift DB to NL. Store all explanations. Specifically, we use the following prompts:
    1. (a)
      Prompt for NL vs SQL views ((i) and (iii)): Which database table or view from the list ¡$R_{2D}$¿ does this phrase ¡r¿ most likely describe? Begin your answer with this table/view.
    2. (b)
      Prompt for NL vs NL views ((ii)): Which database table or view description from the list most likely describes the same table or view as this phrase? Begin your answer with your chosen description from the list.
    3. (c)
      Prompt for NL vs SQL roles ((i) and (iii)): Which database role from the list does this phrase most likely describe?
    4. (d)
      Prompt for NL vs NL roles ((ii)): Which database role description from the list most likely describes the same role as this phrase?
    5. (e)
      SQL vs SQL: simply compare role lists. In the case of views, implement views on DB and determine equality pairwise.
  3. (3)
    For the roles or views that are the same, determine whether the privileges of $M_{2}$ are in violation of those of $M_{1}$ and label them accordingly. Store the explanation. (we prompt ChatGPT as to whether one set of privileges exceeds the other).
LLM4AC At Scale. LLM4AC scalably maps roles and views by using prompts that match single roles/views from $M_{1}$ to lists of roles/views $M_{2}$ (i.e., perform a nearest-neighbor-if-exists on roles/views of $M_{2}$). For example, ChatGPT has a context window length of 8192 tokens, and a list of 10 view definitions, each roughly 50 tokens long, on average, resulting in only 10 calls of 500 tokens each, with 500 more tokens of output. This costs 5500 tokens. This is cheaper compared to 100 pairwise comparisons, each consisting of at least 100 tokens and 1 token of output (a yes/no token) (10001 tokens). This enables cheaper differencing of large NLACMs, such as comparing a policy to a production database.
5. LLM4AC Evaluation
In this section, we propose a benchmark for evaluating IBAC-DB systems (Section 5.1). We use this benchmark to show that our system, LLM4AC, can accurately synthesize and difference access control policies (Sections 5.2, 5.3 and 5.4), especially compared to other baselines. Finally, we show that LLM4AC performs accurately when external knowledge is needed using a real-world access control policy (Section 5.5). For all experiments, the DBMS is PostgreSQL. We use ChatGPT (gpt-3.5-turbo) as the LLM, and ChatGPT’s APIs for all prompting.
5.1. IBAC-DB Benchmark Design
We design a benchmark to test IBAC-DB systems that implement access control policies in PostgresQL. The goal of the benchmark is to test how accurately an IBAC-DB system synthesizes a SQL implementation of a NLACM policy, and differences two NLACMs.
5.1.1. The Differencer Benchmark
Differencer Evaluation Scenarios. We measure differencing with respect to two scenarios: (i) policy auditing: compare a NLACM containing NL to its implementation on the DB, which is a NLACM containing SQL. (ii) policy comparison: compare two NLACMs containing NL.
Differencer Evaluation Metric. We evaluate differencing with respect to F1 scores, where a positive indicates that both NLACMs share a privilege. We also use F1 scores to evaluate role-view mapping and privilege subsumption. A role-view mapping positive indicates a role-view pair is shared by both NLACMs. A privilege subsumption positive indicates the privileges of the second NLACM are subsumed by the first. In cases where performance is near-perfect, we instead use the F1 error (1 - F1) for presentation.
Differencer Benchmark Details. To evaluate LLM4AC with respect to policy auditing and comparison, we construct three NLACMs: a base NLACM containing NL, a SQL NLACM containing the ground truth SQL, and a perturbed NLACM, which is the base NLACM with a NL perturbation applied, such that the resulting NL is syntactically different, but semantically equivalent (e.g., all column names are replaced with synonyms). Then, policy auditing is differencing the base NLACM to the SQL NLACM. Policy comparison is differencing the base and perturbed NLACMs.
An IBAC-DB system needs to have high cross-domain performance–that is, the differencer should accurately compare policies on databases with a variety of schemas and policy specifications. To test this using our benchmark, we construct these NLACMs from: (i) databases spanning multiple training domains, and (ii) syntactically different but semantically equivalent NL (e.g., column names expressed using carrier phrases instead of their actual names). We gather this data from an existing benchmark, Dr. Spider (Chang et al., 2023), a NL2SQL benchmark with databases spanning several training domains, and several NL perturbations.
There are only 5 databases in Dr. Spider that have NL-query-perturbation triples for all perturbation types. We create the equivalent base, SQL, and perturbed NLACMs for each of these databases. We list the databases and some basic characteristics here:
  1. (1)
    orchestra: 4 tables, 27 columns
  2. (2)
    dog_kennels: 8 tables 57 columns
  3. (3)
    employee_hire_evaluation: 4 tables, 21 columns
  4. (4)
[…]
  5. (5)
    car_1: 6 tables, 29 columns
NLACM Construction. To construct NLACM pairs for policy comparison and auditing, we extract the NL describing database views from Dr. Spider. We automatically generate the roles and privileges in the NLACM. To experiment with NL perturbations of roles and privileges as well, we define and use the following perturbations:
  1. (1)
    Role synonyms: replace all role names with synonyms (e.g., “Nonprofit Organization intern” $\rightarrow$ “Charitable Organization Administration Intern”
  2. (2)
    Role descriptions: replace all role names with descriptions of the role. (e.g., “A person who works in a charitable organization to gain experience in overseeing operations and programs.”)
  3. (3)
    Privilege synonyms: replace permitted SQL operators with synonyms. (e.g., “SELECT, UPDATE, INSERT, GRANT” $\rightarrow$ ”This position holds the authority to CHOOSE, ESTABLISH, BESTOW permissions on this database perspective.”)
  4. (4)
    Privilege carrier phrases: replace permitted SQL operators with carrier phrases that imply them. (e.g., “SELECT, UPDATE, INSERT, GRANT” $\rightarrow$ “This role is authorized to perform actions on this database view.”)
The base, SQL, and perturbed NLACMs should be equivalent. That is, the (i,j)th cell of all three NLACMs should describe the same privileges on the same role on the same view. We construct NLACMs that each contain 10 views and 10 roles. We chose this size of NLACM because human users will create NLACMs manually, so it is unlikely that NLACMs will be large.
5.1.2. The Synthesizer Benchmark
Synthesizer Evaluation Scenarios. It is essential that IBAC-DB system synthesizers be able to accurately implement a policy, regardless of the complexity or variety of queries that must be generated. The main difficulty of translating NLACMs to SQL is translating views, as view definitions can be very complex, involving multi-way joins, complex filters, etc. To this end, we measure synthesizing with respect to the types of SQL queries that must be generated when given NL descriptions of database views. We use the following taxonomy:
  1. (1)
    Simple projections (single SELECT over multiple columns)
  2. (2)
    Complex projections (SELECT over multiple columns, aggregates, over multiple conditions of multiple tables, etc.)
  3. (3)
    Simple whole-table aggregations (single GROUP BY over 1-2 attributes in a single table)
  4. (4)
    Simple joins (join between two or three tables on clearly specified FKs)
  5. (5)
    Multi-way joins (join between more than three tables)
  6. (6)
    Common table expressions (uses the HAVING keyword)
  7. (7)
    Simple conditions (1-3 conditions each using single predicates on an attribute, or an aggregate of an attribute)
  8. (8)
    Complex conditions (conditions can be lengthy, there may be self-defined conditions, self-defined variables in the condition, etc.)
Synthesizer Evaluation Metric. We input this NLACM to LLM4AC’s Synthesizer and test the accuracy of the result by executing the resulting GRANT statements on one copy of the databases, and executing the ground truth on a separate copy of the databases. PostgreSQL stores the results as a list of privileges in a table. We compare these tables to see if they are equal.
Synthesizer Benchmark Details. To gather the NL and SQL of these queries of varying types and complexity, we use a mix of the TPC-H queries and queries from several Spider databases.
  • TPC-H: 8 tables. We select queries from the set given in the benchmark, and we extract the NL from the business questions describing each query, available in the TPC-H specification.
  • Spider (Yu et al., 2018): contains 166 separate databases, each containing between 3 and 14 tables. We choose queries from three databases from Spider: department_management, culture_company, and bike_1. The NL is the ground truth input to a NL2SQL system whose ground truth output is the query.
We gather up to 5 queries for each category above from the TPC-H queries and Spider queries, resulting in 33 queries altogether. We modify each query to create a view from its result. Then, we randomly generate 10 roles, and privileges on each view to create one NLACM with dimensions 10 rows and 33 columns.
Figure 6. LLM4AC’s Differencer F1 Scores by Perturbation.
5.2. LLM4AC Differencing: Macrobenchmark
Evaluation Baselines. We compare LLM4AC’s differencing procedure to the following baselines for role-view mapping:
  1. (1)
    Plain LLM: LLM4AC prunes out incorrect answers by only including candidates that share DB literals. We quantify the benefits of literal pruning by including a LLM baseline where answers are not pruned out. Everything else is kept the same.
  2. (2)
    Sentence Similarity: Embed the input role/view and the list of roles/views using SentenceBERT (Reimers and Gurevych, 2019). Then, return the sentence with the least cosine similarity.
  3. (3)
    Token Similarity: Embed the tokens of the input role/view and the tokens of each role/view in the list using BERT. For each role/view in the list, average the token embeddings and choose the role/view whose cosine similarity has the least distance to the average embedding of the input role/view.
  4. (4)
    Jaccard Similarity: Choose the role/view from the list whose tokens have the highest Jaccard similarity with the tokens of the input.
For all baselines except Jaccard, we adjust them to compare NL and SQL by generating a NL description of the SQL query, and comparing it to a list of NL roles/views of the other NLACM.
We do not include baselines for privilege subsumption as it is an ill-defined learning problem and, as we will see, prompting performs sufficiently well on it.
Plain LLM Sentence Word Jaccard LLM4AC
NL vs NL
Total
RV 1-F1
0.149 0.084 0.220 0.204 0.041
NL vs NL
Total
Priv 1-F1
0.119 0.116 0.122 0.116 0.003
NL vs SQL
Total
RV 1-F1
0.553 0.765 0.632 0.837 0.035
NL vs SQL
Total
Priv 1-F1
0.021 0.007 0.014 0.014 0.006
Table 1. LLM4AC Differencer F1 Errors across all DBs and Perturbations. Lower is better.
Table 1 shows total F1 errors calculated over all DBs and perturbations for privilege subsumption and role-view mapping. The first two rows are policy comparison results (NL vs NL), and the second two rows are policy auditing (NL vs SQL).
Policy Comparison.: The first two rows of Table 1 show that, across databases and perturbations, role-view mapping seems more difficult than privilege subsumption. Further, LLM4AC or sentence embeddings outperform other role-view mapping methods.
Policy Auditing.: The last two rows of Table 1 show that overall, LLM4AC outperforms the baselines. For other methods, the F1 errors are all higher compared to those of policy comparison. The F1 error for Plain LLM is generally higher than for other methods, which initially suggests that using LLMs is not preferable. However, this is because all the errors are still due to mistakes that could have been fixed by comparing DB literals. LLM4AC incorporates this observation, leading to its high performance.
5.3. LLM4AC Synthesizer: Macrobenchmark
Query Type Accuracy
Single column Projection 50 / 50
Multiple column Projection 50 / 50
Single Whole-table Aggregation 30 / 30
Single join 50 / 50
Multi-way join 10 / 20
Common Table Expression 40 / 40
Nested Queries 0 / 10
Single WHERE clause condition 50 / 50
Multiple WHERE clause conditions 30 / 30
Total Accuracy 310 / 330
Table 2. LLM4AC Synthesizer Accuracy. The only errors were nested queries and multi-way joins.
Table 2 shows that LLM4AC’s synthesizer is mostly accurate, failing at multi-way joins and nested queries. This is because: (1) C3 can occasionally choose the wrong tables for a multi-way join (2) C3 can “forget” that a nested query is needed.
5.4. LLM4AC Differencer: Microbenchmark
We study the variations in the LLM4AC Differencer performance due to specific NL perturbations or databases. We find that, even after stratifying the results across databases and NL perturbations, LLM4AC still outperforms other baselines (Section 5.4.1). We then study LLM4AC’s performance in-depth (Section 5.4.2).
5.4.1. LLM4AC vs Baselines, by DB/Perturbation
Policy Comparison. When we compute the F1s for each perturbation, we find that LLM4AC vastly outperforms all other methods, followed by sentence embeddings (Figure 6, (A)). This is because naively providing all inputs to prompting generates incorrect answers that are fixed by comparing sentences with shared DB literals instead. On the other hand, sentence embeddings have more obscure errors: namely, when the correct answer is a more abstract sentence (e.g., ”How many orchestras has a conductor done?”) sentence embedding similarity is incorrect. We find the same pattern when stratifying over databases.
Policy Auditing. (Figure 6, (B)) show the results for differencing stratified by perturbation. We do not find a difference between LLM4AC and other role-view mapping methods with respect to perturbation or database.
5.4.2. LLM4AC’s Performance, In-Depth
Figure 7. LLM4AC’s NL vs SQL Role-View Mapping F1s by DB. y-axis is boxplot of F1s across perturbations.
Figure 8. LLM4AC’s NL vs SQL Role-View Mapping F1s by Perturbation. y-axis is boxplot of F1s across databases.
Figure 9. LLM4AC’s NL vs NL Role-View Mapping F1s by DB. y-axis is boxplot of F1s across perturbations.
Figure 10. LLM4AC’s NL vs NL Role-View Mapping F1s by Perturbation. y-axis is boxplot of F1s across databases.
We analyze LLM4AC’s performance variations on various databases and perturbations to better characterize its performance. We do not show LLM4AC’s privilege subsumption performance, as it performs perfectly with no variation across databases and perturbations.
Policy Auditing. In Figure 8, we show LLM4AC’s NL vs SQL F1 scores stratified by perturbation. We find that auditing varies only slightly. Figure 7 shows LLM4AC’s performance stratified by database has almost no variation among different databases. However, LLM4AC’s performance at comparing NL whose column indicators have been replaced with column values is noticeably lower compared to other perturbations. This is because ChatGPT hallucinates by assuming there is no correct answer when matching SQL to NL whose column indicators have been replaced by values in the column. One way to solve this would be to input possible values that can appear in a column as well, as hints to the LLM about which column may have the chosen value.
Policy Comparison. In Figure 10, we show LLM4AC’s NL vs NL F1 scores stratified by perturbation, in order to determine if policy comparison varies by perturbation type. Overall, LLM4AC’s performance shows slight variations by perturbation. Figure 9 shows that LLM4AC’s performance stratified by database has almost no variation among different databases. That said, LLM4AC’s performance at comparing raw NL to its keyword synonym perturbation is noticeably lower compared to other perturbations. This is because ChatGPT hallucinates by assuming the correct answer is not among the answer choices provided (even when there is only one answer choice provided, which is the correct one). This leads to a lower recall, and therefore a lower F1.
5.5. Incorporating External Knowledge
We test LLM4AC on a real-world benchmark that we create using a proprietary database and access control matrix from the Institute for Translational Medicine’s Sociome Data Commons (itm, [n.d.]). We further test LLM4AC on 3 databases of varying sizes from the BIRD NL2SQL benchmark, which also includes external knowledge hints (Li et al., 2023).
The Sociome Benchmark. Each table in the sociome data commons contains columns representing a social, political, or environmental factor (e.g., pollution levels, average household income, etc.). We were also provided with an access control matrix describing a role hierarchy of 6 roles for each of these factors (columns). The roles are ”ALL” (read access for the general public), ”ALL_ACDM” (read access for academic institutions) ”UCHICAGO” (read access only for University of Chicago employees), ”RUSH” (read access only for Rush University employees), ”RESEARCH” (read/write access only for researchers), and ”SOCIOME_INFRA” (full access for the Sociome Data Commons infrastructure team). We constructed 15 views using SQL projections for columns permitted for a role.We construct a 6 role by 15 view NLACM containing the SQL for access control on the Sociome Data Commons by using a provided data dictionary to concatenate the definitions of projected columns and summarize them using GPT-4. We manually generate NL descriptions of the roles. The resulting NL roles and views are the NLACM policy for the sociome database. Lastly, we generate perturbed versions of the policy NLACM by prompting GPT-4 with the NL role/view descriptions, and instructions for perturbing role/view descriptions provided by the Dr. Spider benchmark (Chang et al., 2023). Since we are given only the database schema, (tables and columns), and our view definitions only consist of projections, we only apply perturbations on NL references to columns.
BIRD Benchmark. Each NL-SQL pair has a sentence representing external knowledge. NLACM construction is similar to that of Dr. Spider, but we choose queries deemed ”challenging” or ”moderate” in generation difficulty. We specifically test the impacts of context size on synthesis and differencing performance using 3 databases of low, medium, and high sizes from BIRD: (i) small- california_schools: 3 tables, 89 columns (225 tokens in total), (ii) medium-financial: 8 tables, 55 columns (493 tokens), (iii) large-european_football_2: 201 tables, 1608 columns (1056 tokens). We synthesize and difference 3 10x10 NLACMs, one per database.
Synthesizer Results. The synthesizer correctly generates 54 / 90 on the sociome NLACM when only using the database schema. Adding the data dictionary of column definitions as a rule to C3-AC increased the accuracy to 78 / 90. The BIRD benchmark results were (i) small: 72 / 100, (ii) medium: 70 / 100, (iii) large: 71 / 100. Surprisingly, context size has little impact on performance. Further, we find that errors are specifically due to lack of information about column values. Extending C3-AC rules to incorporate this was sufficient to correct many errors, updating the results to: (i) small: 88 / 100, (ii) medium: 95 / 100, (iii) large: 92 / 100.
Perturbation
NL vs NL
Sociome
RV F1
NL vs NL
Sociome
Priv F1
NL vs SQL
Sociome
RV F1
NL vs SQL
Sociome
Priv F1
column_attribute 0.929 1.0 0.519 1.0
column_value 0.815 1.0 0.462 1.0
column_synonym 0.159 0.8 0.159 1.0
column_carrier 0.846 1.0 0.519 1.0
role_syn 1.0 1.0 1.0 1.0
roledesc_replace 1.0 1.0 1.0 1.0
priv_syn 1.0 0.774 0.519 1.0
priv_inf 1.0 0.774 0.519 1.0
Table 3. Role-View Mapping (RV) and Privilege Subsumption (Priv) F1 Scores on the Sociome Policy
Differencer Results. Table 3 shows the Sociome benchmark’s role-view mapping (RV) and privilege subsumption (Priv) F1 scores when using the data dictionary as input. While performance is high considering the new requirement of external knowledge, errors are due to two key LLM hallucinations: (1) Converting view candidates expressed in SQL to NL when searching for a SQL view that matches a given NL view during role-view mapping (e.g., the LLM response to parse is ”average household income” instead of the column name, ”avg_Agg_HH_INC_ACS”). LLM4AC fails to parse the LLM response to a SQL view in this case. (2) Hallucinating NL view descriptions that were not candidate matches for a given NL view description. We report the NL vs SQL total F1 scores of the BIRD benchmark (there are no perturbations to test NL vs NL, and privilege subsumption is 100% accurate, so there is no need for ablation): (i) small: 0.73, (ii) medium: 0.68, (iii) large: 0.76.
We observe that using column definitions from the data dictionary and column value information substantially boosts performance, and suggests that domain experts can easily boost LLM4AC’s performance for their data domains using data dictionaries through few-shot examples for LLM4AC’s C3-AC and differencing prompts. Even without column value information, LLM4AC’s synthesizer accuracies exceed state-of-the-art NL2SQL scores (bir, [n.d.]). Our high differencing F1 scores are a positive first step.
6. Requirement-Specific Extension
In this section, we explain the design of RHieSys, a system that extends LLM4AC to new access control requirements, namely role hierarchies. The design of RHieSys uses requirement-specific prompting for synthesis and differencing. Specifically, RHieSys works as follows: it takes as input multiple tables, one per requirement type. RHieSys first performs a similarity join over these tables. The resulting table is called an Extended Natural Language Access Control Matrix (ENLACM). Then, RHieSys uses requirement-specific synthesis and differencing procedures on the ENLACM. We describe these steps in detail.
6.1. Using LLMs to audit similarity joins
6.1.1. Approach Overview
The NLACM represents policy for access control privileges, and our goal is to extend it with a policy representation for role hierarchies, the RHL. This requires a similarity join on roles such that the errors will be traceable to pairs of tuples aligned incorrectly. To achieve this, RHieSys performs a similarity join in two passes: given NLACM $M$ and RHL $L$, the first pass performs a similarity join between $M$ and $L$ on $M.role=L.role$, where both $M.role$ and $L.role$ are NL descriptions of roles. The second pass uses $M.role$, $L.role$, and the role pairs aligned by the similarity join of the first pass to audit the join result for potential errors. Note that these two passes can be thought of as role match synthesis and differencing output matches against the original two role sets.
To perform a similarity join, we must entity match access control objects (roles, in this case). This requires solving the following problem: given two sets of objects $O_{1}$ and $O_{2}$ where each access control object $o_{i}\in O_{i}$ describes a database access control object, let $f:O_{1}\times O_{2}\rightarrow\{0,1\}$ be a function such that $f(o_{1},o_{2})=1$ if $o_{1}$ and $o_{2}$ describe the same access control object, and $f(o_{1},o_{2})=0$ otherwise. The goal of entity matching is to find the set of pairs $O_{match}=\{(o_{1},o_{2})\in O_{1}\times O_{2}|f(o_{1},o_{2})=1\}$.
There are many choices of similarity function for the first pass join. Some involve string or set similarity heuristics, leveraging exact matching of strings, thresholding based on edit/Hamming/Jaccard distance, graph matching, etc (a survey of these techniques can be found here (Papadakis et al., 2020)). Others are learned, and involve leveraging various deep learning architectures, such as Siamese networks, RNNs/LSTMs, binary classifiers, and fine-tuned pretrained transformers (a survey can be found here (Li et al., 2021)). Recently, LLMs have been shown to be effective at entity matching on commonly accepted benchmarks as well (Narayan et al., 2022). RHieSys can use any similarity function, but in this paper, we instantiate RHieSys with LLMs with prompts specialized for matching access control objects, due to the lack of representative training data for access control policies, which is most likely to occur in practice.
After performing the similarity join, RHieSys must audit to detect potential errors. That is, access control object pairs aligned by the similarity join $\hat{O_{match}}$ and object sets $O_{1}$ and $O_{2}$, RHieSys must find all $(o_{1},o_{2})\in\hat{O_{match}}$ such that $f(o_{1},o_{2})=0$. In practice, we lack labeled training data to learn $f$, so instead, we solve the problem: find all $(o_{1},o_{2})\in\hat{O_{match}}$ for which there exists $o_{1}^{\prime}\in O_{1}$ or $o_{2}^{\prime}\in O_{2}$ such that $(o_{1}^{\prime},o_{2})$ or $(o_{1},o_{2}^{\prime})$ is more likely to be similar than $(o_{1},o_{2})$. This means a similarity function for auditing must take as input not only $O_{1}$, $O_{2}$, but also the set of object pairs. We use LLMs for auditing, as they generalize most easily to this problem (we can prompt them with all inputs).
6.1.2. Synthesis & Auditing of Similarity Joins
Given $M.role$ and $L.role$, we use a LLM-based bipartite matching procedure for the first pass, first creating a matching between $M.role$ and $L.role$ to use in performing the join. We exploit the insight that every role in $M.role$ is distinct, but can map to one or more roles in $L.role$. For the second pass, given the role matches used to synthesize the join in the first pass, denoted $K$, we audit $K$ for potential errors.
  1. (1)
    Maintain a list of remaining roles in $L.role$, $L^{\prime}$, and a set of pairs $K$ indicating the bipartite matching between $M.role$ and $L.role$.
  2. (2)
    For each role description $r\in M.role$, prompt the LLM as follows: Consider the following role description $r$. List all roles from the list of role descriptions $L^{\prime}$ that describe the same role as $r$. The LLM will output list $O\subset L^{\prime}$. Add pairs $\{(r,o):o\in O\}$ to $K$. Update $L^{\prime}$ to $L^{\prime}\setminus O$.
  3. (3)
    For auditing, maintain a list of errors, $E$. For each pair $(r,o)\in K$, prompt the LLM: Consider the following descriptions of a database role: $r$, $o$. These roles are drawn from a database with roles $M.role$, $L.role$. Do these describe the same role? If the answer is no, store the pair in $E$.
  4. (4)
    Perform an equijoin on $M$ and $L$ using columns $M.role$ and $L.role$, where equality is defined according to $K$. The output is the ENLACM, whose columns will be all columns in $M$ and $L$. Note that this means both columns $L.role$ and $M.role$ are included.
The potential errors $E$ will be returned to the end user, who can correct the resulting ENLACM before using it for synthesis and auditing of access control implementation, which we describe next.
6.2. ENLACM to Implementation
In this section, we describe how RHieSys synthesizes access control implementation from the ENLACM, generates a ENLACM from the database for the access control implementation, and how it differences implementation with the ENLACM policy in order to audit for compliance. Importantly, synthesis must determine when role inheritance GRANT statements should be used instead of granting SQL operations, and differencing must account for this when determining privilege violations. Note that synthesis and differencing are parametrized by the ENLACM in this case, as opposed to the NLACM described in Section 3.
6.2.1. Synthesis Procedure
Given a ENLACM $M_{E}$ and database $B$ with no roles or privileges, Step 1: Generate a ”CREATE ROLE/USER…” statement for each role description from the $M.role$ and $L.role$ columns of $M_{E}$ by prompting the LLM to generate a role label for each role description $r$ using prompt: Consider the following role description $r$. What role does this describe? Store the mapping of role descriptions to labels, denoted $R_{B}$. Step 2: Map each role description $r_{c}$ in $M_{E}.child$ to a label in $R_{B}$ using the following prompt, which takes as input $r_{c}$, the role description $r\in L.role$, its label $l_{r}$, and role labels $R_{B}$: Consider the following role $l_{r}$ described as $r$. $r$ inherits privileges from a role described by $r_{c}$. $r_{c}$ describes one role in the following set: $R_{B}$. What is the best label for this description?. Store the mapping of child descriptions to labels, denoted $R_{C}$. Step 3: Repeat the same process for $M_{E}.parent$. We use an analogous prompt. Store the mapping of parent descriptions to labels in $R_{P}$. Step 4: Generate a ”CREATE VIEW…” statement for each table/view in the ENLACM columns using the following prompt, inspired by C3 (Dong et al., 2023): Complete postgres SQL statement only and with no explanation, and do not grant privileges on tables, roles, and users that are not explicitly requested in the statement. CREATE VIEW. Step 5: For each $M_{E}$ cell in columns representing table/view privileges, given the generated role ¡r¿ and view SQL ¡v¿ and the privilege string in the cell. Generate the GRANT statement for the privilege. We use the following prompt: Consider the following statement: ¡NL for privileges¿. According to this, which of the database operations SELECT, UPDATE/INSERT, DELETE, CREATE, GRANT are permitted for role ¡r¿ on table/view ¡v¿? Let $G_{M}$ be the set of output matrix of GRANT statements, such that the $(i,j)$th cell of $G_{M}$ is implements the set of privileges for the $(i,j)$th cell of $M_{E}$. Step 6-Implementing Inheritance: For each GRANT statement in $G_{M}$, incorporate inheritance if needed by recursively finding the children’s GRANT statements, parsing them for allowed SQL operators, taking a union of the operators, and then replacing the original GRANT statement with statements ”GRANT $r_{c}$ TO $r$”, and a GRANT statement for operators not inherited.
6.2.2. From SQL to ENLACMs for Auditing
To audit an implementation for compliance, RHieSys has a Generator that parses the GRANT statements for both privileges and inheritance in the database (stored in a metadata table specifying privileges, such as postgres’s information_schema. role_table_grants) to generate a ENLACM containing the SQL of the policy. The Generator constructs the ENLACM in two passes: (i) parse GRANT statements that grant permissions on SQL operators to ENLACM cells. (ii) Parse role inheritance GRANT statements and replace each role’s permitted SQL operators with the union of their SQL operators and their children’s SQL operators.
6.2.3. Differencing Procedure for Auditing
Given 2 ENLACMs $M_{1}$ and $M_{2}$, where $M_{1}$ contains the NL policy, and $M_{2}$ contains the SQL implementation, the goal of differencing is to ensure (i) that $M_{2}$ does not contain privileges not described in $M_{1}$, and (ii) that $M_{1}$ and $M_{2}$ contain the same role inheritance relationships. Therefore, the differencing procedure returns: (i) roles and views in $M_{2}$ not in $M_{1}$, (ii) privileges in $M_{2}$ that are more permissive than $M_{1}$, (iii) role-child pairs in $M_{1}$ but not in $M_{2}$ (iv) role-child pairs in $M_{2}$, but not in $M_{1}$.
Notation. Given NLACM $M_{i}$, let $R_{i}$ and $V_{i}$ be the set of roles and views of $M_{i}$ respectively, and let $C_{i}$ be the set of role-child relationships. We classify the roles/views of each of $M_{1}$ and $M_{2}$ as NL or SQL based on whether the text contains SQL keywords or not. Let $R_{iD},R_{iN}\subset R_{i}$ be the roles of $R_{i}$ containing SQL and the roles of $R_{i}$ only in NL, respectively. We define view sets $V_{iD}$ and $V_{iN}$ and child sets $C_{iD}$ and $C_{iN}$ similarly.
Then, the full procedure is:
  1. (1)
    For both $M_{1}$ and $M_{2}$, classify roles/views as NL or SQL.
  2. (2)
    For each role/view string of $M_{1}$, use heuristics to find substrings likely to be DB literals (e.g., strings representing numbers, strings in quotes, etc.), and prune out candidate roles/views from $M_{2}$ with mismatching literals. The output will be dictionaries $\{r\in R_{1}:R^{\prime}_{2}\subset R_{2}\}$ and $\{v\in V_{1}:V^{\prime}_{2}\subset R_{2}\}$, where $R^{\prime}_{2}$ and $V^{\prime}_{2}$ are the $M_{2}$ roles and views that share literals with role $r$ and view $v$ from $M_{1}$.
  3. (3)
    Determine which roles and views are the same, and which are different using a prompt for NL roles/views to NL roles/views, or a NL to SQL prompt. Store all explanations. Specifically, we use the following prompts:
    1. (a)
      Prompt for NL vs SQL views: Which database table or view from the list ¡$V^{\prime}_{2D}$¿ does this phrase ¡$r$¿ most likely describe? Begin your answer with this table/view.
    2. (b)
      Prompt for NL vs NL views: Which database table or view description from the list ¡$V^{\prime}_{2N}$¿ most likely describes the same table or view as this phrase? Begin your answer with your chosen description from the list.
    3. (c)
      Prompt for NL vs SQL roles: Which database role from the list ¡$R^{\prime}_{2D}$¿ does this phrase most likely describe?
    4. (d)
      Prompt for NL vs NL roles: Which role description from the list ¡$R^{\prime}_{2N}$¿ most likely describes the same role as this phrase?
    5. (e)
      SQL vs SQL: simply compare role lists. In the case of views, implement views on DB and determine equality pairwise.
  4. (4)
    For the roles or views that are the same, determine whether the privileges of $M_{2}$ do not comply with those of $M_{1}$ and label them accordingly. Store the explanation. (we prompt the LLM as to whether one set of privileges exceeds the other).
  5. (5)
    For the roles that are the same, determine which role-child pairs are the same and which are different using prompts similar to those for finding shared roles and views. Output role-child relationships in $M_{1}$ but not in $M_{2}$, and vice versa.
6.3. Integrating Outputs
We cannot naively apply both synthesis and differencing procedures for privileges and role hierarchies, because the correct implementation depends on inheritance vs privileges. In this section, we describe our procedure for combining synthesis and differencing outputs using the semantic join mapping and a prompting strategy to resolve extra privileges that may have been assigned along with inheritance.
7. RHieSys Evaluation
The previous evaluation (see Section 5) evaluated the performance of synthesis and differencing purely with respect to access control privileges. Our goal is to evaluate an EIBAC-DB system with respect to role hierarchy structure. We create a benchmark to evaluate any EIBAC-DB system for role hierarchies and privileges using employee-manager relationships from the Amazon Access dataset, a real-world dataset. We find that RHieSys accurately performs similarity joins, synthesis, and differencing irrespective of role hierarchy tree structure. Further, we systematically test RHieSys using various types of equivalent NL to express roles and privileges, and still find that RHieSys performs well at synthesizing and differencing role hierarchies. For all experiments, RHieSys uses ChatGPT (gpt-3.5-turbo) and PostgreSQL as the database backend.
7.1. Benchmark Design
To test joins and synthesis, we must generate NLACM-RHL pairs containing NL to represent the access control policy. To test differencing, we must generate two inputs: one NLACM-RHL pair, and one ENLACM containing the SQL implementation. Further, role hierarchies can have different structures, and so we must generate multiple NLACM-RHL pairs for different role hierarchies.
Role Hierarchy Structures The Amazon Access dataset has 888 different anonymized role hierarchies, of variable sizes (depths, widths). Notably, most trees are small (only 2 nodes). We ignore such trees and use 3 different trees: deep, a 16-node tree with 6 levels, 12 leaves. wide, 19-node tree with 2 levels, 1 root, and 18 leaves, and balanced, a 12-node tree with minimum 3 levels, maximum 4 levels.
RHL Generation Because the roles are anonymized, we must generate hierarchical role descriptions ourselves. To achieve this, we prompt GPT-4 using the following procedure: (i) we start from the root node which has role, ”CEO”, and we prompt GPT-4 for a role that would be directly below this position in a corporate role hierarchy. (ii) We use a breadth-first traversal. For all non-leaf nodes, we prompt GPT-4 (gpt-4-turbo) for a manager role below the parent node label. (iii) For leaf nodes, we prompt GPT-4 for an employee role below the parent node label. We reorganize the resulting NL tree into a RHL with columns role, parent, and child. We also store the tree of role descriptions, denoted $T_{R}$, for use in generating NLACM privileges, as we will describe.
[…]
Implementation Generation To generate ENLACMs representing implementation, we manually create role labels that would be used in a GRANT statement for each RHL role description. We create a new RHL containing these role labels. We then create the NLACM as follows: the role column of the RHL is the role column of the NLACM, the tables/views are the ground truth SQL for the NL queries chosen from Dr. Spider, and the cells are the lists of permitted SQL operators. We obtain the ENLACM through a plain equijoin of RHL and the NLACM on the role column.
Role and Privilege Perturbations There can be multiple ways of describing the same role or set of privileges. To ensure RHieSys is robust to this, we systematically perturb the role and privilege NL for all NLACM-RHL pairs using the Dr. Spider-inspired perturbations for roles and privileges from Section 5.
Evaluation Metrics We evaluate similarity joins by the frequency of correct tuples that appear in the join. To achieve this, we use F1 scores, where a positive indicates that a NLACM role is equivalent to a RHL role. We evaluate synthesis as the number of access control privileges and inheritance statements correctly implemented. To achieve this, we execute each statement synthesized by RHieSys and its ground truth on separate copies of the database. PostgreSQL stores the privileges and inheritances in its metadata tables. We compare equivalence of recordsets from both copies of metadata tables to determine whether the synthesized statement is correct. We evaluate differencing of policy ENLACM $M_{1}$ against implementation ENLACM $M_{2}$ using F1 scores, where a positive indicates that $M_{1}$ and $M_{2}$ share a privilege or inheritance. A negative indicates that an inheritance is not shared by $M_{1}$ or $M_{2}$, or $M_{2}$ defines a privilege not defined in $M_{1}$.
7.2. Similarity Join Results
We test RHieSys’s similarity join performance by role structure and role perturbation. The results are shown in Table 4.
wide deep balanced
base 0.92 0.86 0.96
role_syn 1.00 0.97 0.97
roledesc_replace 0.97 0.97 1.00
Table 4. Similarity Join F1 Scores, by role structure (left to right) and role perturbation (top to bottom).
In general, RHieSys performs well at similarity joins. We do not find conclusive performance differences with respect to role structure. Errors are primarily due to false positives from allowing ChatGPT to choose multiple candidate role matches from the RHL (see Section 6.2). This is likely due to LLMs’ bias for generating or selecting verbose answers, as observed in recent studies (Saito et al., 2023; Huang et al., 2024; Dai et al., 2024).
7.3. Synthesizer Results
We test RHieSys’s synthesizer performance by role structure and role/privilege perturbations. The results are shown in Table 5.
wide deep balanced
base 0.82 0.82 0.84
role_syn 0.79 0.79 0.75
roledesc_replace 0.80 0.81 0.77
priv_syn 0.83 0.84 0.88
priv_inf 0.86 0.87 0.92
Table 5. Synthesizer Accuracy, by role structure (left to right) and role perturbation (top to bottom).
We do not find conclusive performance differences with respect to role structure. However, we find that RHieSys consistently performs slightly better with respect to the priv_inf and priv_syn perturbations compared to others. This is because base has slightly fewer role inheritance errors compared to role perturbations (e.g., 1 / 18 incorrect inheritance relationships, compared to 2 / 18 on roledesc_replace, for the wide role structure).
In general, we observe that synthesis errors are due to three sources: (i) incorrectly choosing the child role label, leading to incorrect inheritance (maximum 3 / 19 errors observed) (ii) incorrect GRANT statements for allowed SQL operators, due to incorrect inheritance (maximum 44 / 190 errors observed) (iii) incorrect generation of view SQL (maximum 1 / 10 errors observed) due to string column value hallucination.
7.4. Differencer Results
We test RHieSys’s differencer performance by role structure and role/privilege perturbations. The results are shown in Table 6.
wide deep balanced
base 0.82 0.78 0.80
role_syn 0.82 0.74 0.85
roledesc_replace 0.83 0.77 0.82
priv_syn 0.87 0.81 0.86
priv_inf 0.88 0.84 0.86
Table 6. Differencing F1 Scores, by role structure (left to right) and role perturbation (top to bottom).
Performance is consistent across role/privilege perturbations and role structures. We observe that errors are due to (i) incorrect differencing of views (maximum of 3 out of 10 errors observed across perturbations and role structures). When comparing natural language view descriptions to SQL defining views, ChatGPT has the general tendency to choose no views from the list (e.g., when prompted to select from a list containing a single answer, which is correct, ChatGPT will still determine that no view description matches the given SQL). (ii) Ignoring mismatching views causes RHieSys to also ignore implemented privileges for the view that actually are subsumed by the policy. The majority of errors are due to this. (iii) We also observe a small number of errors from incorrectly determining whether role relationships are shared. When determining if relationships are equivalent, ChatGPT tends to overlook differences between positions that seem similar, such as ”Lead Software developer” and ”Software developer”.
8. DePLOI: An Initial System
8.1. Requirement Identification
Given policies $P$, policy documentation $P_{Doc}$, and database backend $B$, the goal is to provide a phrase summarizing the type of deployment requirement captured by these policies. This may involve extracting information about the type of deployment requirement from documentation, if it was provided by the user.
LLMs are known to excel at information extraction and text classification tasks, so we use the following simple prompt: Consider the following table: $P_{i}$. Consider the given documentation describing this table: $P_{Doci}$. What category of deployment information does this represent?
8.2. Dependency Resolver
The dependency resolver takes as input backend $B$, policy tables $P$, policy documentation $P_{Doc}$, and requirement labels generated during requirement identification $L_{R}$. The output is an execution graph $G=(V,E)$ where nodes are policy tables and edges represent required inputs. We implement and test the dependency resolver in two ways: (i) using a prompt that generates the instructions $Inst$, (ii) using human instructions. We use Chain-of-Thought prompting as the prompt for dependency resolver, following the intuition that we are really attempting to generate instructions for implementing deployment requirements on $B$ using the NL in $P$ and $P_{Doc}$: You are a database security analyst. Your job is to make sure excessive access privileges are not granted to anyone. Consider the following tables representing deployment policies: $L_{R}$, $P$. Consider the following documentation for these tables: $P_{Doc}$. Write a set of instructions for implementing these policies on a $B$ backend. Once we have instructions, then we ask followup questions to generate $G$: For each step in the instructions, which policy tables from $L_{R}$ are needed? and For each step in the instructions, list the previous steps whose outputs will be needed for this step.
We construct $G$ such that: each node $v\in V$ is a list of tables, and $e\in E$ exists between nodes $v_{1}$ and $v_{2}$ if the outputs of instruction $v_{1}$ are needed for $v_{2}$.
8.3. Primitive Implementation
Using the given requirement labels and execution graph, DePLOI uses a set of general and extensible primitives to synthesize and audit deployment policy implementation. DePLOI achieves this using the following templated prompts for synthesis and auditing.
8.3.1. Synthesis Primitives
We use the following primitive:
synthesize($P$, $P_{Doc}$, $C$, $B$, $D$, $G$) which outputs database code in the language used by backend $B$.
This primitive is implemented using the following prompt: Implement the following deployment constraints in database backend $B$: Requirements: $P$ Documentation Describing the Tables: $P_{Doc}$, Instructions for implementation: $G$. Also make sure to use the following schema information, where applicable: Database Schema: $D$, Schema information: $C$. Generate the required SQL statements.
We parse the LLM output string for all code and execute this against the database.
8.3.2. Auditing Primitives
First, we must generate implementation tables $I$ for each deployment requirement implementation. This is done in two steps: (i) implementation extraction: extract the elements of the database state (e.g., database privileges) used to implement deployment requirements into a set of implementation tables, one for each deployment requirement. An implementation table’s semantics are the same as those of policy. For example, an access control privilege policy table’s columns, rows, and cells represent tables/views, roles/users, and privileges, respectively, and an implementation table’s columns, rows, and cells will represent the same types of information. However, in an implementation, the cells, column headers, and row indices consist of code rather than NL. (ii) dependency-based revisions: simply generating the implementation tables from database state is not enough–the semantic relationships among implementation tables of different requirements must match those of policy tables. For example, an access control privilege table may contain all allowed SQL operators, both inherited and non-inherited. However, the implementation table will only consist of GRANT statements containing non-inherited operators. Therefore, generating a cell of the access control privilege implementation table, which describes all allowed operators on the database for a role on a table/view would require a union over all inherited and non-inherited privileges. Currently, we observe that LLMs struggle to discover such complex relationships and correctly revise the implementation tables to reflect them using only policy tables and their documentation. Therefore, we assume users give us a revision function to apply on the implementation tables, $Rev(I)$.
Implementation Extraction. Interestingly, we empirically observe that LLMs lack knowledge required to generate code that will query the database for access control configuration details, e.g., GRANT statements and TRIGGERs. Therefore, rather than generating code using LLMs, we simply allow users to input the queries needed to extract the implementations of various deployment requirements (e.g., current access control privileges on Postgres, current temporal access conditions, etc.). For example, for an access control privilege policy table, the generator will output a table containing code representing policy implementation, rather than NL describing policy. The $(i,j)th$ cell of an implementation table for access control privilege implementation would contain ”GRANT SELECT ON customer TO ceo;”, where the row represents all database privileges granted to ”ceo” and the column represents all database privileges granted on the ”customer” table.
Dependency-based Revisions. The function $Rev(I)$ will revise implementation tables $I$ so that the semantic relationships of $I$ match those of $P$. For example, for access control requirements including temporal access control, access control privileges, and role hierarchies, $Rev(I)$ will: (i) use the role hierarchy table to union privileges assigned to every role with all inherited privileges and (ii) For each cell of the access control privilege table, replace the code in the cell by merging any code granting temporal privileges on a role/view from the temporal access control table with code granting non-temporal privileges on the same role/view.
Differencing Once such implementation tables $I$ have been generated, they must be differenced to ensure they comply with policy $P$. Given policy table and corresponding implementation table $P_{i}$ and $I_{i}$, we determine which privileges are in implementation that are not permitted according to policy in two steps: (i) deployment element identification DePLOI must first extract the types of database elements needed to define a given type of deployment requirement (we call this set of types a deployment element). For example, in a table describing role hierarchies with a ”Role” and ”Child” column, the deployment element is an inheritance relationship, consisting of both parent and child. (ii) deployment element matching: determine which database elements (e.g., roles/users, tables/views) appear in both $P_{i}$ and $I_{i}$ (iii) privilege comparison: for database element matches, determine whether implementation privileges in $I_{i}$ are permitted according to privileges in $P_{i}$. If not, then $I_{i}$ does not comply with $P_{i}$, and DePLOI records the mismatching privilege.
8.4. Synthesis Feedback
We observe that LLMs generate queries with correct syntax, but incorrect column values or misunderstanding foreign key relationships. A general approach used by many LLM-backed NL2SQL systems to solve such problems is to carefully select the correct database context to include as input to the LLM such that the LLM will output the correct database query (Floratou et al., 2024). This context includes both database schema elements $D$ and database values. Context selection involves determining which database elements are likely to appear in the ground truth database query for a NL description. It is difficult due to NL ambiguities, but also due to the sheer scale of most databases. We observe that rather than attempting to select this context as input, a more successful strategy is to pass in a base amount of context with the NL initially, and then parse the output queries using existing DBMS parsers (e.g., using sqlglot, or mongo-parse) for the tables, columns, and column values they represent, check the parsed elements against the database for non-existent tables, columns and column values, and then use the discovered errors as feedback to the LLM to regenerate a more accurate query.
9. DePLOI Evaluation
In this section, we discuss how to evaluate a system for synthesizing and auditing formal specification such as DePLOI. We develop an initial macrobenchmark and detailed microbenchmark for evaluating DePLOI with respect to synthesis and auditing accuracy. Then, we provide a detailed design for a more in-depth evaluation of DePLOI.
Evaluation Challenges. As mentioned in Section 1, DePLOI will be used to reduce the human effort required in developing formal specifications from policy documents themselves. As a first step toward evaluating whether DePLOI reduces human effort, we view system accuracy as a proxy for human effort, as the fewer errors a human has to detect and correct, the less effort they will require. Further, since DePLOI is the first generalized automated system to synthesize and audit deployment policy, there is no existing system to evaluate DePLOI against in a macrobenchmark. As a first step toward a macrobenchmark, we compare the system accuracy of DePLOI against a task-specific variant. We then discuss a full design for a macrobenchmark involving a user study.
We evaluate DePLOI on a realistic benchmark and find that DePLOI performs almost as well as a task-specific alternatives, and that DePLOI’s performance is not affected by choice of input and metadata formats. However, we also observe that DePLOI’s performance is limited due to current limitations of LLMs, giving rise to novel research challenges in using LLMs for deployment.
9.1. Benchmark Details
We evaluate DePLOI with respect to access control deployment requirements: access control privileges, role hierarchies, and temporal access control. This means the benchmark requires policy tables consisting of: (i) a role hierarchy (ii) Descriptions of database views (iii) descriptions of allowed SQL operators (iv) descriptions of temporal conditions, e.g., ”9 to 5 on weekdays” (v) the ground truth implementations for each of the above on all backends we will test. We describe them in detail here.
Policy Tables. There are 3 policy tables: (i) a natural language access control matrix (NLACM) consisting of columns whose headers are NL descriptions of database views, a ’Role’ column containing NL descriptions of roles, and cells describing permitted database operations for a role on a database view. (ii) a role hierarchy list (RHL) consisting of a ’Role’ and ’Child’ column, where each row represents a role that inherits privileges from a child role. Each cell contains a NL description of a database role. (iii) a temporal access control matrix (TACM), with a ’Role’ column and columns representing database views, similar to a NLACM, but whose cells represent temporal constraints.
We construct each of these tables using real-world datasets and state-of-the-art NL2SQL benchmarks.
Generating Roles & Views We use database views derived from NL questions representing SQL queries from the recent BIRD benchmark. These queries are over the european_football_2 database, which is the database with the largest schema in the benchmark, with 8 tables and 201 columns. Note that we use the database with the largest schema (as opposed to, say, database size) because schema size affects the size of context needed as input to DePLOI, which affects performance. We use roles from a real-world role hierarchy of the Amazon Access dataset (cite amazon access). As the roles in this dataset are anonymized, we use GPT-4 to generate NL descriptions of roles for a corporate role hierarchy. These generated role and view descriptions are used in the cells of the ’Role’ column and as the column headers of the NLACM and TACM, and we use the hierarchy tree given by the Amazon Access dataset to organize the generated role descriptions into ’Role’ and ’Child’ columns in the RHL.
NLACM Cell Generation For each NLACM cell, we randomly generate permitted database operations while respecting the role hierarchy (that is, for roles that inherit privileges, we only randomly generate additional non-inherited privileges).
TACM Cell Generation For each TACM cell, we randomly generate temporal filters by generating components of periodic expressions, which have been used to describe temporal access control in prior work (cite trbac).
Policy Table Documentation We provide column definitions and relationships between policy tables, similarly to documentation provided with open tables, such as the UC Irvine Machine Learning Repository and government open data portals.
Implementation Tables for Differencing To test auditing (DePLOI’s differencing procedure), the benchmark also must include implementation tables containing the ground truth database deployment code for each policy table. We use the ground truth SQL from the BIRD benchmark for database views, the role names matching each description for database roles, and we manually generate implementation for the access control privileges, role hierarchy, and temporal access control conditions.
Unless otherwise stated, we use ChatGPT as the LLM, Postgres as the database backend, and JSON objects as the format for metadata and schema information.
9.2. Macrobenchmark: DePLOI vs Task-Specific
We compare DePLOI against a task-specific version of DePLOI, which uses prompts and synthesis and auditing procedures specific for implementing access control privileges, temporal access control, and role hierarchies.
Synthesis. DePLOI correctly generates 194 / 285. The task-specific variant correctly generates 243 / 285. However, DePLOI has fewer failure modes, which may make it easier for the end user to debug. We analyze the errors more closely: DePLOI has view errors due to deployment element mismatches reintroduced by the feedback loop (38), incorrect privilege-inheritance generation (12), and incorrect temporal constraint generation due to stray GRANT statements (41). On the other hand, the task-specific variant has the same failure modes, and more: view errors due to PK/FK misunderstanding or column values (19), incorrect privilege-inheritance generation (3), incorrect temporal constraint generation due to stray GRANT statements (5), or due to incorrect temporal constraint specification (15).
Auditing. DePLOI’s novel differencing procedure has a F1 score of 0.71. The task-specific variant has a F1 score of 0.85. DePLOI’s sources of error are false positive role inheritance matches, false negative temporal constraint matches and minor feedback-related LLM hallucinations The task-specific variant’s sources of error are false positive role inheritance matches.
9.3. Microbenchmarks
The reference architecture for DePLOI has a large design space characterized by LLM choice, database backends, and schema and metadata formats. We evaluate DePLOI with respect to each of these design space aspects, and we find that many LLMs simply lack the knowledge to generate deployment requirement implementations.
9.3.1. LLM Choice
We evaluate DePLOIusing closed, small open, and large specialized open LLMs. We use gpt-3.5-turbo (ChatGPT) as the closed LLM, mistral-7b-instructv0.1 as the small open LLM, and CodeLlama-70B as the large specialized open LLM.
Synthesis. ChatGPT has an accuracy of 194 / 285 due to generation of code that does not conform to the given policy. On the other hand, Mistral-7b has an accuracy of 127 / 285 due to incorrect syntax–parts of queries are missing. This is partially because Mistral-7b has a small context window compared to ChatGPT, meaning its responses are often incomplete. CodeLlama-70B has an accuracy of 165 / 285. While CodeLlama shares the same failure modes as ChatGPT, we observe that CodeLlama also often hallucinates by assuming it is not possible to generate the correct implementation without more information.
Auditing. ChatGPT has an F1 score of 0.71. On the other hand, Mistral-7b has an F1 score of 0.49 because it very often hallucinates reasoning on privilege subsumption and role inheritance in various ways (e.g., reasoning that UPDATE is less permissive than SELECT, or that two role inheritance relationships with identical strings for the roles represent different relationships, etc.). CodeLlama-70B has an F1 score of 0.61 due to similar hallucinations in reasoning about privilege subsumption and role inheritance. However, these hallucinations are fewer.
9.3.2. Database Backend
We evaluate DePLOI with respect to two backends with different SQL variants for access control implementation (Postgres and MySQL) and one non-SQL backend (MongoDB).
Synthesis. DePLOI on Postgres has an accuracy of 194 / 285. On MySQL, the accuracy was 180 / 285–the results are exactly the same, but for one temporal condition, ChatGPT does not realize it cannot use a CREATE POLICY statement, as MySQL does not support row-level security directly. On the other hand, MongoDB has an accuracy of 42 / 285. The high number of errors is due to major python parsing issues and errors, and ChatGPT is completely unable to implement temporal constraints.
Auditing. Postgres has an F1 score of 0.71, Auditing for MySQL has an F1 score of 0.64. The errors are exactly the same as for Postgres, but there are further errors due to ChatGPT not recognizing that some TRIGGER statements match to temporal conditions. Auditing for MongoDB is 0.12. Errors are due to many hallucinations where ChatGPT refuses to determine whether python programs conform to a specification, or simply responds a correct temporal constraint does not match its NL specification.
In summary, DePLOI’s performance depends on the LLM’s knowledge of database backend semantics, but LLMs’ knowledge of certain backends, like MongoDB, is currently limited.
9.3.3. Schema Format
We evaluate DePLOI with respect to three different formats for representing the database schema and data dictionary metadata from the catalog to ChatGPT: as JSON objects where keys are table names and values are columns, as tables with a ’Table’ column and a ’Column’ column, and as text documents.
Synthesis. DePLOI on JSON achieves accuracy 194 / 285. DePLOI on Tabular achieves accuracy 189 / 285. Here, the results are exactly the same, except for 5 hallucinations arbitrary hallucinations. DePLOI on Unstructured Text achieves accuracy 191 / 285. Here also, the results are exactly the same, except for 3 arbitrary hallucinations–distinct from the hallucinations of Tabular.
Auditing. DePLOI on JSON achieves an F1 score of 0.71. DePLOI on Tabular achieves an F1 score of 0.65. Here, the results are exactly the same, except for 9 arbitrary hallucinations. DePLOI on Unstructured Text achieves 0.68. Here, the results are exactly the same, except for 4 arbitrary hallucinations.
In summary, we observe that varying the format in which database schemas and data dictionaries from catalogs are shown to the LLM has little effect on DePLOI’s performance.
9.3.4. Policy Format
We evaluate DePLOI with respect to two different formats for representing the policy: the tabular representations we have proposed, such as NLACMs, and documents consisting of sentences where each sentence represents a table cell, describing either permitted SQL operations, temporal constraints, or a role hierarchy relationship.
Synthesis. DePLOI on tables achieves accuracy 194 / 285. DePLOI on documents achieves accuracy 187 / 285. Here, the results are exactly the same, except for 7 extra GRANT statements.
Auditing. DePLOI on tables achieves an F1 score of 0.71. DePLOI on documents achieves an F1 score of 0.65. Here, the results are exactly the same, except for 9 arbitrary hallucinations.
In summary, we observe that varying the format in which deployment policies shown to the LLM has little effect on DePLOI’s performance. While documents may underperform slightly compared to tables, we conjecture this only occurs because representing documents as a string to LLMs requires more tokens than tables, and LLMs currently cannot perform accurately with large volumes of context. The difference in performance with respect to policy formats will likely diminish as LLMs improve.
9.3.5. Can LLMs Handle Dependency Resolving?
DePLOI assumes instructions for executing code are given as documentation by the user. However, DePLOI would reduce synthesis and auditing effort even more if LLMs did not need such documentation to resolve code dependencies among different deployment requirements. We experiment with few-shot and Chain-of-Thought prompts on the more powerful GPT-4 to generate the output execution graphs, but we find that DePLOI severely underperforms because GPT-4 cannot reason about such dependencies. We report the highest performance we observed over various prompting strategies. The best strategy was Chain-of-Thought.
Synthesis. DePLOI with LLM-based dependency resolving achieves 0.432 compared to 0.681 using user-provided documentation. This is because in addition to correct CREATE POLICY statements to implement temporal constraints, DePLOI synthesizes incorrect GRANT statements which permit the same SQL operations without the required temporal constraints. Despite providing GPT-4 several few-shot examples of this reasoning, GPT-4 underperformed.
Auditing. DePLOI with LLM-based dependency resolving achieves 0.53 compared to 0.71 using user-provided documentation. Similarly to the failure mode for synthesis, this is because in addition to correct CREATE POLICY statements to implement temporal constraints, DePLOI falsely indicates errors due to missing GRANT statements which would incorrectly permit the same SQL operations without the required temporal constraints.
10. Conclusion
In this paper, we recognize the problem of automating the policy comparison and auditing of access control policies written in NL. To facilitate this, we propose IBAC-DB, a new access control paradigm. We define a reference architecture for IBAC-DB which can be applied across several different LLMs and DBMSs. Then, we build one possible system, LLM4AC, using this reference architecture. We find that LLM4AC performs well compared to other role-view mapping methods, and that LLM4AC performs reliably: it has low performance variation by database and NL perturbation. We build systems to handle other database deployment requirements as well, RHieSys and DePLOI, and we find promising initial results.
Appendix A DePLOI: Human Inputs
DePLOI differs from RHieSys and LLM4AC in that it uses human inputs, namely the dependency resolver instructions and the policy documentation. We provide the exact text of these instructions and documentation here.
A.1. Dependency Resolver Instructions
We show our dependency resolver instructions in Figure 11. Note that these instructions are agnostic to the input policy format.
Figure 11. Human Instructions given as input to DePLOI
A.2. Policy Documentation
The policy documentation describes the format in which the policy is shown, the deployment requirements represented, and their semantics. We provide the exact text of this documentation provided as input to DePLOI in Figure 12.
Figure 12. Policy Documentation given as input to DePLOI. Documentation for the table policy format on top, document policy format on bottom.